Protiern??7SNI A 5nT7 znick layer of oil wth densi
Answer #1

gauge pressure at bottom of container = Patm + (rho1 * g * h1) + (rho2 * g * h2)

                                                            = (1.013 * 105) + (0.05 * 9.8 * 920) + (1000 * 9.8 * 0.1)

                                                            = 1.0273 * 105 Pa